# Approximating the coefficients of the Bessel functions

### v4

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
over $B_{R}(0)$ by Lemma˜10.27, applying Corollary˜10.17 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$. Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$. If $\eta$ is compactly supported, then the free convolution is compactly supported and therefore uniquely determined by its moments by Carleman’s condition, which implies that the sequence of probability measures also converges weakly to the free convolution after applying [billingsley]*Theorem 30.2.
If $\alpha=0$, then the convergence in terms of moments is clear and the weak convergence follows from [billingsley]*Theorem 30.2. ∎
@@ 11. Computing the Dunkl bilinear form with combinatorics
The previous result is not applicable to when $k+\ell(\lambda)-\ell(\nu)\geq N$, but if this is the case, then we must consider when the number of distinct indices is $N$. Then, to obtain a summation formula for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ in (A), we must set $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ to be one of the summands. Similarly, we cannot obtain analogous formulas for (B) and (C).
@@
Assuming that $\theta$ is fixed, the order of the expressions in Theorem˜11.1 matches the order of the expressions in Theorems˜6.1, 7.1 and 8.10, which is $N^{k+\ell(\lambda)-\ell(\nu)}$. However, the leading order coefficients are not apparent from the formulas.
12. Leading order terms of the type A and BC Dunkl bilinear forms after multiplying by $x_{1}\cdots x_{N}$
In this section, we prove analogues of Theorems˜8.16 and 8.23. First, we prove the analogues of Theorem˜8.16 for the $A^{N-1}$ and $BC^{N}$ root systems.
Theorem 12.1.
$[ep_{\lambda},ep_{\nu}]_{A^{N}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(N\theta)^{2k-\ell(\nu)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}+R(N,\theta)\bigg),$
where $R\in\mathbb{Q}[x,y]$ satisfies:
  1. (1)
  2. (2)
    The degree $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the degree of $y$ is at most $2k-\ell(\nu)$.
Proof.
The same method as the proof of Theorem˜8.16 can be used. ∎
Corollary 12.2.
Suppose $\lambda,\nu\in\Gamma$ and $|\lambda|=|\nu|$. For (B), assume that $\lambda,\nu\in\Gamma_{\text{even}}$. Suppose $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
$(A)\,\,\,[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+(i-1)\theta)[p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$
$(B)\,\,\,[ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}=(1+o_{N}(1))\prod_{i=1}^{N}(1+2(i-1)\theta)[p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$
Proof.
(A) follows from Theorems˜6.1 and 12.1 and (B) follows from Theorems˜8.10 and 8.16. ∎
Theorem 12.3.
$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})\bigg(N^{2k+\ell(\lambda)-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}(2\theta_{0})^{k-\ell(\nu)}\times$
$\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}+R(N,\theta_{0},\theta_{1})\bigg),$
where $R\in\mathbb{Q}[x,y,z]$ satisfies:
  1. (1)
  2. (2)
    The degree of $x$ is at most $2k+\ell(\lambda)-\ell(\nu)-1$ and the sum of the degrees of $y$ and $z$ in each summand is at most $2k-\ell(\nu)$.
  3. (3)
    In no summand is the degree of $x$ $\ell(\lambda)$ greater than the degree of $y$ while the degrees of $y$ and $z$ add to $2k-\ell(\nu)$.
Proof.
First, we have that $h\mathcal{D}(e)ep_{\nu}=\mathcal{D}(he)hep_{\nu}=\mathcal{D}(e)ep_{\nu}$ for all $h\in H(BC^{N}(\theta_{0},\theta_{1}))$. Hence, $\mathcal{D}(e)ep_{\nu}$ has all even degrees. It follows that we can replace $\mathcal{D}(ep_{\lambda})$ with
$\prod_{i=1}^{\ell(\lambda)}\sum_{j=1}^{N}\mathcal{D}_{j}^{\lambda_{i}-1}\partial_{j}\mathcal{D}(e),$
as we have done previously.
After noting that $\theta_{1}$ has order $\theta_{0}N$, the same method as the proof of Theorem˜8.16 can be used. In this case, we can view the switch from $i$ to $j$ as
$\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right).$
We replace $\frac{\theta_{1}(1-\tau_{i})}{(N-1)x_{i}}$ with $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}$ to compute the leading order term. Thus, we can view the switch from $i$ to $j$ as
(39) $\frac{\theta_{1}(1-\tau_{i})}{Nx_{i}}+\theta_{0}\left(\frac{1-s_{ij}}{x_{i}-x_{j}}+\frac{1-\tau_{i}\tau_{j}s_{ij}}{x_{i}+x_{j}}\right);$
afterwards, we follow the same method. The goal is to compute
$\sum_{s\in\mathcal{S},\,r\in\mathcal{R}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}s_{|\lambda|}\circ\cdots\circ s_{1}\circ r_{k-d_{1}(s)}\circ\cdots\circ r_{1}\prod_{j=1}^{k-d_{1}(s)}x_{j}p_{\nu},$
where $\mathcal{S}$ is defined in the same way except that the switch from $i$ to $j$ is replaced with (39) and $\mathcal{R}(s)$ is defined in the same way except that $2\theta d_{i}$ is replaced by $(2\theta_{0}+\frac{2\theta_{1}}{N})d_{i}$. We get that we can similarly reduce to the case that $\ell(\lambda)=1$. Hence, let $c_{\nu^{\prime}}$ be the value of (22) when $\nu$ is set as $\nu^{\prime}$ and $\lambda$ is set as $(|\nu^{\prime}|)$ for $\nu^{\prime}\in\Gamma_{\text{even}}$.
The only difference is in Lemma˜8.20. Assume that $\pi\in NC^{\prime}(|\nu^{\prime}|)$ and $b(\pi)=\nu^{\prime}$. For $s\in\mathcal{S}$ and $r\in\mathcal{R}(s)$ that are associated with $\pi$, we have that the number of switches in $s$ which are not associated with a derivative in $r$ is $|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1$ and the number of switches in $r$ is $|\nu^{\prime}|-\ell(\nu^{\prime})+1$; recall Definition˜8.17 for the definition of $d_{3}(r)$. The switches in $r$ contribute a factor of $\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}$. On the other hand, the switches in $s$ which are not associated with a derivative in $r$ contribute a factor of $(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}(1+\frac{\theta_{1}}{\theta_{0}N})^{o(\pi)}$. The switches in $s$ which are associated with a derivative in $r$ contribute a factor of $(-2\theta_{0})^{d_{3}(r)}$. Hence, the total contribution is
$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{d_{1}(s)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-\ell(\nu^{\prime})+1}(2\theta_{0})^{|\nu^{\prime}|-d_{1}(s)-d_{3}(r)-1}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}(-2\theta_{0})^{d_{3}(r)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-d_{3}(r)}(-1)^{d_{3}(r)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)+d_{3}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)},$
where we have used $d_{1}(s)+d_{2}(r)+d_{3}(r)=\ell(\nu^{\prime})-1$. Observe that the number of locations for a switch in $s$ that is associated with a derivative in $r$ is $z(\pi)$. For each of these locations, we can decide to place the derivative in $s$ or in $r$. After summing over these choices, we obtain
$\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-d_{2}(r)-z(\pi)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}-2\theta_{0}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})+d_{2}(r)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-d_{2}(r)}$
$=\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
Thus, we obtain that
$c_{\nu^{\prime}}=$ $\pi(\nu^{\prime})\prod_{l=1}^{\ell(\nu^{\prime})}\nu^{\prime}_{l}\sum_{\pi\in NC^{\prime}(|\nu^{\prime}|),\,b(\pi)=\nu^{\prime}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{|\nu^{\prime}|-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}(2\theta_{0})^{|\nu^{\prime}|-\ell(\nu^{\prime})}\times$
$\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}.$
It follows that
$\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}c_{\gamma((\nu_{j}:j\in S_{i}))}$
$=\prod_{l=1}^{\ell(\nu)}\nu_{l}\sum_{[\ell(\nu)]=S_{1}\sqcup\cdots\sqcup S_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{\lambda_{i}-z(\pi)}\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\times$
$(2\theta_{0})^{\lambda_{i}-\ell(b(\pi))}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC^{\prime}(\lambda_{i}),\\ b(\pi)=\gamma_{i}\end{subarray}}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}$
$=\pi(\nu)\prod_{l=1}^{\ell(\nu)}\nu_{l}(2\theta_{0})^{k-\ell(\nu)}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{k}\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC^{\prime}(\lambda_{i})}\left(2\theta_{0}+\frac{2\theta_{1}}{N}\right)^{-z(\pi)}\times$
$\left(\frac{2\theta_{1}}{N}\right)^{z(\pi)}\left(1+\frac{\theta_{1}}{\theta_{0}N}\right)^{o(\pi)-\mathbf{1}\{f(\pi)>1\}}\prod_{B\in b(\pi)}x_{|B|}.$
After multiplying by $\prod_{i=1}^{N-k}(1+2(i-1)\theta_{0}+\theta_{1})N^{2k+\ell(\lambda)-\ell(\nu)}$, we obtain the desired result. ∎
Next, we prove the analogues of Theorem˜8.23 for the $A^{N-1}$ and $BC^{N}$ root systems.
Theorem 12.4.
$[ep_{\lambda},ep_{\nu}]_{A^{N-1}(\theta)}=\prod_{i=1}^{N-k}(1+(i-1)\theta)\bigg($ $N^{\ell(\lambda)}(1+N\theta)^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{D;\,\text{odd}}(\pi)(N\theta)\prod_{B\in\pi}x_{|B|}$
$+R(N,\theta)\bigg),$
Proof.
We can prove this in the same way as we prove Theorem˜8.23. ∎
For $\pi\in NC(k+1)$, define
$W^{BC;\,\text{odd}}(\pi)(x,y)\triangleq$ $(1+\mathbf{1}\{b(2;\,\pi)=0\}(d(2;\,\pi)-1))\times$
$\prod_{i\in[k+1],i\geq 3,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi))\left(\frac{1+y}{1+x+y}\right)^{z(\pi)}\times$
$\left(\frac{1}{1+x+y}\right)^{\mathbf{1}\{f(\pi)>1\}}.$
Theorem 12.5.
$[ep_{\lambda},ep_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $\prod_{i=1}^{N-k}(1+2\theta_{1}+2(i-1)\theta_{0})\times$
$\bigg(N^{\ell(\lambda)}(1+2\theta_{1}+2N\theta_{0})^{k}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\times$
$\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i}+1)}W^{BC;\text{odd}}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in b(\pi)}x_{|B|}$
$+R(N,\theta_{0},\theta_{1})\bigg),$
Proof.
We can prove this theorem in the same way as we prove Theorem˜8.23. The main difference is that in the last step of the proof of Lemma˜8.24, the total contribution is
$1-\frac{2N\theta_{0}}{1+2\theta_{1}+2N\theta_{0}}=\frac{1+2\theta_{1}}{1+2\theta_{1}+2N\theta_{0}},$
which would correspond to $\frac{1+y}{1+x+y}$ rather than $\frac{1}{1+x+y}$. ∎

### v5

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
over $B_{R}(0)$ by Lemma˜10.29, applying Corollary˜10.18 again gives that the $k$th free cumulant of $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$. Hence, the probability measures $\mu^{\theta}_{\alpha N}(\eta_{N})(x\mapsto\frac{x}{\theta N})$ converges to the free convolution of $\eta$ and the semicircle law multiplied by $\frac{\alpha}{\theta}$ in terms of power sums as $N\rightarrow\infty$. ∎
Example 10.31.
Next, we study the law of large numbers for the $\beta$-Laguerre ensemble. First, we compute the type BC Bessel generating function of the Chiral ensemble, which has the same distribution as the square roots of the $\beta$-Laguerre ensemble, where $\beta=2\theta$. We discuss its construction for $\theta\in\{\frac{1}{2},1,2\}$ and its definition for general values of $\theta\geq 0$ from [forrester]*Section 3.1. Suppose $M\geq N$ such that $M-N+1-\frac{1}{2\theta}\geq 0$. Let $X$ denote a random $M\times N$ matrix whose independent entries are real, complex, or real quaternion numbers with densities:
  • $\frac{1}{\sqrt{2\pi}}e^{-\frac{x^{2}}{2}}$ if $\theta=\frac{1}{2}$,
  • $\frac{1}{\pi}e^{-|x|^{2}}$ if $\theta=1$,
  • $\frac{4}{\pi^{2}}e^{-2|x|^{2}-2|y|^{2}}$ if $\theta=2$. Note that we parameterize a quaternion as $(x,y)\in\mathbb{C}^{2}$.
The Chiral ensemble is defined as the positive eigenvalues of
$H=\begin{bmatrix}0_{M\times M}&X\\ X^{H}&0_{N\times N}\end{bmatrix}$.
This matrix almost surely has $N$ negative eigenvalues, MNM-N zero eigenvalues, and $N$ positive eigenvalues.
Define the probability distribution $C_{\theta,M,t}^{N}$ over $\mathbb{R}_{\geq 0}^{N}$ to have the density of $a$ proportional to
$\prod_{i=1}^{N}a_{i}^{2\theta(M-N+1)-1}e^{-\frac{a_{i}^{2}}{2t}}\prod_{1\leq i<j\leq N}|a_{i}^{2}-a_{j}^{2}|^{2\theta}.$
Then, $C_{\theta,M,t}^{N}$ is the density function of the positive eigenvalues of $H$ after they are rescaled when $\theta\in\{\frac{1}{2},1,2\}$. The following lemma is a straightforward generalization of [rectangularmatrix]*Proposition 5.21.
Lemma 10.32.
Suppose $\theta\geq 0$, $M\geq N$ such that $M-N+1-\frac{1}{2\theta}\geq 0$, and $t>0$. Then,
$G_{C^{N}_{\theta,M,t}}^{BC^{N}(\theta,\theta(M-N+1)-\frac{1}{2})}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right).$
Next, we discuss the $\beta$-Laguerre ensemble, which has the same distribution as the squares of the eigenvalues of the Chiral ensemble. Suppose $M\geq N$ such that $M-N+1-\frac{1}{\theta}\geq 0$. Define the probability distribution $L_{\theta,M,t}^{N}$ over $\mathbb{R}^{N}_{\geq 0}$ to have the density of $a$ proportional to
$\prod_{i=1}^{N}a_{i}^{\theta(M-N+1)-1}e^{-\frac{a_{i}}{2t}}\prod_{1\leq i<j\leq N}|a_{i}-a_{j}|^{\theta}.$
Although we cannot generalize the random matrix that we previously defined for $\theta\in\{\frac{1}{2},1,2\}$, [betaensembles] computes a tridiagonal matrix whose eigenvalue distribution is given by $L_{\theta,M,t}^{N}$ for all $\theta>0$. Moreover, the distribution is the image of $C_{\theta,M,t}^{N}$ under the pushforward map $x\mapsto x^{2}$. Despite this relation, it remains challenging to compute the Bessel generating function for the $\beta$-Laguerre ensemble.
Afterwards, we obtain Corollary˜10.33, which has been previously proved in Theorem 1.1 of the paper [betalaguerre]. The paper proves that the $\beta$-Laguerre ensemble weakly converges to the Marchenko-Pastur law. It is straightforward to verify that the moments appearing in the corollary are the same as those of the Marchenko-Pastur law.
Corollary 10.33.
Assume that $\lim_{N\rightarrow\infty}\theta N=\infty$, $M\geq N$, $M-N+1-\frac{1}{2\theta}\geq 0$, $\lim_{N\rightarrow\infty}\frac{M}{N}=c\geq 1$, and $\lim_{N\rightarrow\infty}t=\alpha\geq 0$. Then, for all $\nu\in\Gamma$,
$\lim_{N\rightarrow\infty}\frac{\mathbb{E}_{a\sim L^{N}_{\theta,M,t}}[p_{\nu}(a)]}{(\theta N)^{|\nu|}N^{\ell(\nu)}}=\prod_{i=1}^{\ell(\nu)}\sum_{p\in\mathcal{D}_{\nu_{i}}}c^{e(p)}(2\alpha)^{\nu_{i}}.$
Proof.
Use Corollary˜10.19, Lemma˜10.32, and the fact that $L_{\theta,M,t}^{N}$ is the image of $C_{\theta,M,t}^{N}$ under the pushforward map $x\mapsto x^{2}$. ∎
@@ 11. Computing the Dunkl bilinear form with combinatorics
The previous result is not applicable when $k+\ell(\lambda)-\ell(\nu)\geq N$, but if this is the case, then we must consider when the number of distinct indices is $N$. Then, to obtain a summation formula for $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ in (A), we must set $p_{\lambda},p_{\nu}]_{A^{N-1}(\theta)}$ to be one of the summands. Similarly, we cannot obtain analogous formulas for (B) and (C).
@@
Assuming that $\theta$ is fixed, the order of the expressions in Theorem˜11.1 matches the order of the expressions in Theorems˜6.1 and 7.1, which is $N^{k+\ell(\lambda)-\ell(\nu)}$. However, the leading order coefficients are not apparent from the formulas.